OpenStudy (anonymous):

what is 0.3939393939 in a fraction

OpenStudy (anonymous):

can you use the calculator? if so there is a button that calculates this if not; I would round to nearest hundreth (0.39) the decimal needs to be moved 2 places to the right. so I would divide by 100

OpenStudy (anonymous):

What button is it?

OpenStudy (anonymous):

Please help?

OpenStudy (anonymous):

which calc are you using?

OpenStudy (anonymous):

the button typically will look like "F < > D." An F and a D will each have arrows pointing at it.

OpenStudy (anonymous):

I don't have it :( I only have a four function calculator

OpenStudy (anonymous):

Help!!!!!!!!!!!!!!

OpenStudy (anonymous):

NP..then do the other method 39/100

OpenStudy (anonymous):

.39

OpenStudy (anonymous):

100x = 39.3939 - x = 0.393939 -------------- 99x = 39.

OpenStudy (anonymous):

Now, solving 99x=39 for x by dividing both sides of it by 99, we'll get that x=39/99. And this is your answer.

OpenStudy (anonymous):

ohhh thank you!!!!!!1 I got it :)

OpenStudy (anonymous):

no leave it in a fraction that is your question. Here is an online calculator that does it for you also http://www.webmath.com/dec2fract.html

OpenStudy (anonymous):

no my teacher wants it to be in a fraction

OpenStudy (anonymous):

39/99 is a fraction

OpenStudy (anonymous):

.39 is a decimal

OpenStudy (anonymous):

No I am not calling you stupid...you asked me before was the answer .39 I was telling you for that question - no leave it in a fraction..

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

NP...Happy studies